### 17 Mar 2026 v2

# AI4EOSC: a Federated Cloud Platform for Artificial Intelligence in Scientific Research

Abstract
In this paper, we describe the architecture and implementation of a federated compute platform dedicated to support Artificial Intelligence and Machine Learning in scientific workloads. Putting the effort into reproducible deployments, it delivers consistent, transparent access to a federation of physically distributed e-Infrastructures, with a particular focus on the European Open Science Cloud. Through a comprehensive service catalogue, the platform is able to offer an integrated user experience covering the full Machine Learning lifecycle, including model development (with dedicated interactive development environments), training (with GPU resources, annotation tools, experiment tracking, and federated learning support) and deployment (covering a wide range of deployment options all along the Cloud Continuum). The platform also provides tools for traceability and reproducibility of AI models, integrates with different Artificial Intelligence model providers, datasets and storage resources, allowing users to interact with the broader Machine Learning ecosystem. Finally, it is easily customizable to lower the adoption barrier by external communities.
keywords:
Cloud Platform , Artificial Intelligence , Machine Learning , European Open Science Cloud , MLOps
journal: Future Gener. Comput. Syst.firstfirstfootnotetext: First and second authors contributed equally to this work. Other authors are listed in alphabetical order.
[…]
[ifca]organization=Instituto de Física de Cantabria (IFCA), CSIC-UC, addressline=Avda. los Castros s/n, city=Santander, postcode=39006, state=Cantabria, country=Spain
[…]
[upv]organization=Instituto de Instrumentación para Imagen Molecular (I3M), Centro Mixto CSIC - Universitat Politècnica de València (UPV), addressline=Camino de Vera s/n, city=Valencia, postcode=46022, state=Valencia, country=Spain
[…]
[iisas]organization=Institute of Informatics, Slovak Academy of Sciences (IISAS), addressline=Dúbravská cesta 9, city=Bratislava, postcode=84507, country=Slovakia
[…]
[stuba]organization=Faculty of Informatics and Information Technologies, Slovak University of Technology, addressline=Ilkovičova 2, city=Bratislava, postcode=84216, country=Slovakia
[…]
[infn]organization=Istituto Nazionale di Fisica Nucleare (INFN), addressline=Via Enrico Fermi 40, city=Frascati, postcode=00044, state=Roma, country=Italy
[…]
[predictia]organization=Predictia Intelligent Data Solutions, addressline=Fernando de los Ríos 48, city=Santander, postcode=39006, state=Cantabria, country=Spain
[…]
[kit]organization=Karlsruher Institut für Technologie, addressline=Kaiserstraße 12, city=Karlsruhe, postcode=76131, country=Germany
[…]
[lip]organization=Laboratório de Instrumentação e Física Experimental de Partículas, addressline=Av. Prof. Gama Pinto 2, city=Lisboa, postcode=1649-003, country=Portugal
[…]
[pnsc]organization=Poznańskie Centrum Superkomputerowo Sieciowe, addressline=Jana Pawła II 10, city=Poznan, postcode=61-139, country=Poland
[…]
[cnca]organization=Centro Nacional de Computação Avançada (CNCA), addressline=Avenida do Brasil, 101, city=Lisboa, postcode=1700-066, country=Portugal
1 Introduction
Artificial Intelligence (AI), Machine Learning (ML), and Deep Learning (DL) are transforming scientific and industrial research. The impact of these techniques over the last years has revolutionized the field in a wide range of applications—e.g., by making it possible to tackle new scientific endeavors [21, 48, 31], by automatizing manual and tedious tasks [51] or by speeding up some time-consuming and and effort-intensive tasks [85], just to cite some examples. These techniques are a reality, and they form nowadays part of the basic set of tools for a wide set of scientists and researchers [42].
However, this data-driven research revolution demands more than just access to computational and storage resources. Scientists willing to effectively exploit these state-of-the-art techniques require advanced platforms and tools that can lower adoption barriers (e.g., access to accelerators), easing and simplifying data exploration and analysis, model development, training, sharing and deployment into production environments; in an open, transparent and collaborative way. All these aspects go beyond simple model development and require addressing complex engineering challenges that span the entire AI/ML development lifecycle.
In a partial response to these challenges, Machine Learning Operations (MLOps) has emerged as a set of practices and tools designed to ease the management of the entire AI/ML lifecycle [9], analogous to how DevOps bridges the gap between software development and IT operations. MLOps focuses on model deployment automation, reproducibility, and scalability, facilitating the seamless transition from data science experimentation to the reliable operation of ML components in production. While MLOps has been successfully adopted by industry, its adoption in the scientific and research context is not that prominent. Many existing MLOps platforms lack important open-science-centric features that scientists require nowadays, such as seamless integration with research infrastructures and data repositories, or direct support for FAIR principles [91].
This gap is particularly critical in the context of the European Open Science Cloud (EOSC) ecosystem. The EOSC is an European Commission initiative aimed at creating the “web of FAIR data and services”, as a transformative process towards the promotion of open science practices among researchers, also delivering a comprehensive environment for the sharing, processing, and analysis of research data [24]. In the context of AI/ML applications for scientific research, realizing this vision requires not only infrastructure but also platforms that can operationalize open science principles across the entire AI/ML lifecycle.
In this context, we present the AI4EOSC platform, a federated system that provides a comprehensive cloud-based toolbox for the development, training, sharing and deployment of AI/ML solutions, with a strong focus on traceability, reproducibility, and open science principles. AI4EOSC enables data processing from the cloud-to-edge continuum, being able to mobilize different resources for different computational needs. Moreover, being developed in the context of the EOSC ecosystem, the platform delivers seamless technical and semantical interoperability, ensuring and strengthening open science practices as a collaborative, transparent, and accessible research model.
The AI4EOSC platform is not merely focused on providing scientists with the tools and functionalities needed to develop, train or serve models, but also to frame the development of these models with the open science practices —and hence, the FAIR principles [91]— by design, ensuring fine-grained model provenance metadata generation our automated metadata interoperability across different formats. To this aim, we have implemented these functionalities in an automated and transparent way for the users, so that these principles can be granted by design, to the possible extent, without their intervention.
AI4EOSC operationalizes and adopts MLOps in a research context, extending its principles to support the unique needs of open science. The platform provides scientists with the robust, reproducible, and scalable workflows they need, ensuring compliance with FAIR principles and comprehensive integration with the EOSC and European Research Infrastructures ecosystem.
The remainder of this manuscript is structured as follows: first, Section 2 introduces the background and related work. In Section 3 and Section 4, we discuss the general architecture of the AI4EOSC platform and we introduce further considerations that underpin the global design and architecture respectively. In Section 5 we explain how different scientists with multidisciplinary profiles can leverage the platform to carry out different tasks. Finally, Section 6 draws out the main conclusions and the future work proposed.
2 Related work and motivation
While there are many private and proprietary platforms to deploy AI/ML workloads for different kind of users, few of them have put an effort on addressing the open science and reproducibility requirements that are key to scientific workloads. There is a clear framework fragmentation, where the existing solutions address a specific part of the AI/ML lifecycle, often being isolated. Moreover, we argue that there is a clear lack of integrated and open-science focused solutions that span the whole AI/ML lifecycle.
2.1 Existing solutions: a fragmented ecosystem
2.1.1 Model development, training and reproducibility
The foundational phase in AI/ML research is the model development and training. This process is normally supported by a wide range of frameworks and tools like Scikit-learn, JAX, PyTorch or TensorFlow [63], providing the needed functionality for model creation, with a strong focus on performance and scalability. These tools normally require access to accelerators (e.g., GPU, TPU, NPU) in order to be able to efficiently build an AI/ML model.
Experiment tracking platforms, including open source —e.g., MLflow [93] or LabML [45]— and proprietary ones —like Weights and Biases [25], Valohai [88] or Comet [23]— allow developers to monitor and compare training runs, but their focus remains on industry workflows, with limited support for open science practices (due to the lack of structured metadata in some cases) or the lack of integration with research infrastructures. Similarly, reproducibility platforms like Code Ocean  [^1] https://codeocean.com/ or Binder [49] allow the creation of reproducible environments, but these are typically isolated from the broader AI/ML lifecycle, lacking connections to model sharing or deployment.
There are some AI/ML platforms that do cover the whole lifecycle process end-to-end. However, these are closed source (e.g., H2O.ai [41]) and restrict model deployment or the data storage to their specific cloud solutions (like Google Vertex AI [4], Amazon Sage Maker [6] or Azure ML Studio [58]), thus defeating the open principles on which science is built. Others are partially open-source but do not offer all features in their community free versions (Polyaxon [72]).
Beyond these commercial solutions, the DEEP-Hybrid-DataCloud project proposed a distributed architecture to provide ML developers with a set of cloud services that allowed them to ease the model creation, training, validation, and model serving [56]. The DEEP platform focused on easing the access to accelerators (GPUs) in pan-European e-Infrastructures, but it lacked technical or semantical interoperability with other systems, especially due to the recent developments and advancements in the EOSC federation and the European Research Infrastructures.
2.1.2 Model sharing and model repositories
A number of AI model repositories exist, including those domain-specific that are devoted to scientific communities, like the BioImage Model Zoo [66] for microscopy images or Kipoi [7] for genomics. Also, community-specific model repositories like BioModels [39, 57] are being augmented, in order to include AI/ML assets along with mathematical models of biological and biomedical systems. Although these systems implement and follow the FAIR principles to different extents, there is a lack of model and metadata interoperability, as well as a lack of integration with broader MLOps techniques.
Besides domain-specific systems, there are large-scale general-purpose model repositories, like OpenML [11] (an open platform to share datasets and AI/ML models), Hugging Face [92], TensorFlow Hub (now integrated in Kaggle) or Replicate [74]. Although these repositories are focused on sharing, model metadata is often unstructured and limited, and most model repositories are detached from the training process of the model, thus making users unable to assess the provenance or trustworthiness of a given asset.
2.1.3 Model deployment and MLOps
The deployment of AI/ML models into production services is a critical step to move ahead from research into real-world applications. However, the current landscape of MLOps tools and practices [9] is very fragmented, with most solutions designed for industry needs, rather than research.
Online AI/ML model serving can be implemented by leveraging a large set of lightweight solutions, allowing users to deploy a model behind an API. General-purpose serving can be directly implemented via Fast API [73], which is widely used as a framework to build REST APIs. However, these solutions are normally developed ad-hoc and lack interoperability. DEEPaaS API [37] allows users to offer a common interface for AI/ML models by exposing their Python code via an HTTP endpoint. TensorFlow Serving allows to serve TensorFlow models via gRPC and HTTP endpoints. BentoML allows users to package the models along with their dependencies into a deployable and self-contained asset. High-performance serving and inference can be implemented via more specific systems, like the NVIDIA Triton Inference Server (supporting multiple AI/ML frameworks and model formats, including TensorFlow, ONNX or PyTorch). More specifically, vLLM [53] has emerged as a specialized and optimized tool to serve large language models (LLMs).
MLOps platforms such as Kubeflow [52] or Seldon Core [83] provide an industry-grade and robust framework to automate model deployment and monitoring, including continuous integration and delivery (CI/CD). However, this strong focus on industry-driven workflows overlooks the unique needs of scientific research, including interoperability, reproducibility, provenance tracking, and FAIR compliance. For example, Kubeflow supports scalable deployments on Kubernetes, but it lacks interoperability with other cloud management frameworks or the possibility to easily execute a given workload on a different cluster technology (like in High Performance Computing systems).
2.1.4 Open science, FAIR principles
The FAIR principles have become one of the main pillars of open-science practices, with significant investments in methods, tools and best practices in order to broader their adoption. As a matter of fact, the FAIR principles underpin the evolution of the EOSC, which, as already stated, aims to facilitate the “Web of FAIR Data and Services for Science”, incentivizing also broader open science practices by improving the research assessment system [24].
In the context of AI/ML research, these aspects should not be considered just as a mere set of best practices, but as an important vehicle towards AI/ML model transparency, reproducibility and trustworthiness. The FAIR principles, when applied to data leveraging platforms like Zenodo or Dataverse, can address data-related risk factors for AI (such as data provenance or licensing) [36]. However, these platforms are not able to respond to the dynamic nature of AI/ML model development and deployment, hence limiting the FAIR principles adoption beyond data. Researchers that want to ensure the FAIRness of their complete AI/ML solutions must bundle together a set of tools, standards and practices, at the cost of reproducibility, efficiency of their research, or compliance with open science principles.
2.2 AI4EOSC motivation
In this fragmented context, there is an opportunity to implement an open-source solution that can address the AI/ML lifecycle end-to-end, from development to deployment, following the open science and FAIR principles by design, therefore giving a more comprehensive overview of all the processes involved, thus fostering reproducibility. That platform should be approachable enough for non-technical scientists to use and offer a ready-to-use environment and workflows to streamline AI model creation. The resulting models should be reusable and not tied to the platform that created them. Besides, in the current ecosystem of European Research Infrastructures and the EOSC Federation, such a platform must be federated and dynamic in nature, allowing to integrate a wide variety of resources, from disparate providers.
The AI4EOSC platform addresses these challenges through its federated architecture, providing a seamless platform to help scientists to develop, train and deploy AI models in a user-friendly manner, following the open science and the FAIR principles by design. AI4EOSC operationalizes open science for AI/ML research, providing a unified, end-to-end platform, that implements open science principles to modern, data-driven research workflows.
The platform presented in this work, whose architecture will be described in Section 3, is an evolution of the DEEP platform [56]. This platform consisted on three different systems, all connected to a shared set of storage and data services: The DEEP training facility, that users could leverage to train their models; the DEEP Open Catalog, a marketplace where uses could share the models; and the DEEP as a service, allowing to deploy ML models on a serverless architecture and query them though a standardized model API.
Building on the basis of the former DEEP platform, AI4EOSC aims to solve the specific challenges that have been identified in this section (among others), through the whole AI/ML lifecycle:
  • 1.
    Research-first and FAIR-by-design platform, prioritizing reproducibility, collaboration, and transparency, with rich metadata for datasets, models, and workflows, ensuring that all research outputs are findable, accessible, interoperable, and reusable without placing additional burdens on researchers.
  • 2.
    MLOps practices tailored for research, extending MLOps principles to support reproducibility, provenance, and FAIR compliance by design via platform-level and user-defined CI/CD pipelines, integrated with data and model monitoring tools.
  • 3.
    End-to-End provenance metadata for AI/ML models, capturing the full lineage of AI/ML processes, from data preprocessing to model deployment, including resource utilization. AI4EOSC leverages standardized provenance models (W3C PROV), ensuring that every step is traceable and reproducible.
  • 4.
    Semantic (via MLDCAT-AP) and technical interoperability with other platforms, such as AI4Life, ensuring seamless interactions with other marketplaces, making it possible to implement and extend these integration towards other technologies.
  • 5.
    Seamless EOSC Integration: AI4EOSC is natively integrated with the EOSC, enabling researchers to leverage EOSC’s infrastructure for data sharing (e.g., EOSC EU Node resources), collaboration, and long-term preservation. This integration ensures that AI/ML research is not only FAIR but also interoperable with the broader European research ecosystem.
  • 6.
    Unified development environment with extensibility and flexible customization methods, allowing to easily integrate additional tools, like image annotation or federated learning.
  • 7.
    Federated workload management system, being able to scale across many physically distributed datacenters and e-Infrastructures, with a variety of deployment and execution methods.
3 The AI4EOSC platform architecture
As already introduced, the AI4EOSC Platform is designed to help scientists to develop, train and deploy AI models in a user-friendly manner. By allowing transparent access to a large set of federated computing resources, the platform ensures scalability and efficiency of scientific workflows. In the following section, we describe the architecture of AI4EOSC, using the C4 Model [13] to provide a clear and structured visualization of its components and interactions.
Our architecture follows a modular design, organized into four distinct systems: AI4EOSC Development Platform, AI4EOSC AI/ML as a Service, AI4EOSC LLM services, and AI4EOSC Platform Orchestration. These components work interconnected in order to be able to deliver an specific functionality, but being integrated with the rest of the systems. This modularity is illustrated in Figure 1. The legend for all the C4 diagrams is shown in Figure 2
Figure 1: AI4EOSC architecture system landscape view (C4 notation).
Figure 2: Legend for all C4 diagrams used in this manuscript.
3.1 AI4EOSC Development Platform system
This system, depicted in Figure 3, includes all the needed services to deliver a comprehensive platform and toolbox to enable researchers to build, develop, train and share AI/ML models following the FAIR principles. It also includes some core, common and shared services that conform the main AI4EOSC control plane, that serve as auxiliary services to both this system and others. The most relevant components are explained next subsections.
Figure 3: AI4EOSC Development platform architecture: container view (C4 notation).
3.1.1 Workload Management System
The Workload Management System (WMS) is a federation of service providers that is used to manage the platform resources and execute the platform jobs. Given the federated nature of the platform, we have chosen a set of components that allow us to federate geographically distributed resources easily in the form of a service mesh. Our WMS comprises the usage of Consul [43] to federate the providers (following a WAN approach) and Nomad [44] to run the workloads. The current federation includes resources from over 5 different clouds located in 5 different countries: IFCA-CSIC (Spain), IISAS (Slovakia), INCD (Portugal), Tubitak (Turkey) and PNSC (Poland).
The WMS runs both user-specific jobs, and system/platform tasks that are executed independently of the users.
User jobs include both AI modules (3.1.3) and tools (3.1.4). AI modules can run in two modes: standard and batch. In standard mode, resources are allocated to users without a predefined time limit, and users access them via an Interactive Development Environment (IDE) or an API. In batch mode, resources are allocated to users for the duration strictly needed to run the training process and then freed for other users.
In addition, we have enriched the Nomad jobs with a number of sidecar tasks to enhance the user experience transparently, like connecting the job with external storage systems, downloading and syncing datasets from external repositories at launch time or delivering user notifications.
System jobs include, apart from the sidecar tasks already mentioned, Traefik [54] reverse proxies to allow users to access their containers, the deployment of the Platform API and the Platform Dashboard (3.1.2), and other internal management and house-keeping tasks that are never exposed to the users.
3.1.2 Platform API and dashboard
The Platform API (PAPI) is the component that delivers the platform functionality via a well-defined set of endpoints, following a REST architecture. It is developed following the OpenAPI specifications, providing the formal description and documentation of the PAPI functionality.
Built on top of the API, the AI4EOSC Dashboard is the main user entrypoint to the AI4EOSC Platform. Both PAPI and the dashboard are built with extensibility in mind, in order to be able to easily integrate new functionality without interfering with the current one.
Figure 4: A snapshot of the AI4EOSC dashboard (available at https://dashboard.cloud.ai4eosc.eu) providing an overview of the existing AI modules in the platform.
The Dashboard (Figure 4 allows users to easily explore and view the associated metadata for the different catalogs available in the system: AI modules (3.1.3), tools (3.1.4), LLM services (3.3) and other external AI catalogs (4.2.2). It also provides the ability to launch development environments or codespaces in the platform, batch executions, easy management of existing deployments, platform-wide secret management (e.g., API keys, storage integration) and resource consumption reporting, among other user-facing functionalities.
3.1.3 AI Modules
The AI modules are the main platform asset that users interact with, as they encapsulate the AI/ML models that the platform is able to manage. They are both produced and consumed by the users, although the AI4EOSC development team also maintains reference implementations of relevant AI models. They are created and stored as git repositories (currently hosted on GitHub [^2] https://github.com/ai4os-hub, but any other software forge can be used) allowing users to easily manage and upload new updates to the modules, as well as making it possible to enforce quality assurance mechanisms and MLOps practices (3.1.5) from the platform point of view. To support the development of new AI modules that comply with the AI4EOSC platform and best practices (e.g., metadata standards (4.1)), a set of pre-defined software templates is provided to users.
3.1.4 Tools
Tools are a collection of additional assets that the AI4EOSC Platform provides to users in order to support them during the whole ML development lifecycle. These tools profit from the platform extensibility (4.2) as a means to enhance and augment the current functionalities.
Unlike AI modules (3.1.3), tools are developed and curated by the developers of the AI4EOSC Platform, in order to ensure that the integration is implemented correctly. Once integrated, tools are easily self-deployable by users on their personal workspaces, just like AI modules.
For technical reasons, some of these tools do not make sense to be deployed per-user and therefore are instead offered as platform-wide services, like the experiment tracking service (3.1.6) or the drift monitoring (3.1.6); however the integration mechanism remains the same.
In the following, we will explain the main set of tools offered by the platform.
Development Environment – Codespaces
The AI4EOSC Platform allows users to deploy a ready-to-use development environment, referred to as Codespaces, to create their new AI modules (3.1.3). The environment can be used out-of-the-box with multiple versions of the main Deep Learning frameworks such as TensorFlow [1], PyTorch [70].
Additional frameworks are also supported by starting from provided plain Ubuntu or NVIDIA CUDA flavours, or by installing a custom software stack.
When configuring the deployment, users can choose their preferred IDE for code development. The platform currently supports both VSCode [59] (bundled with many useful extensions, and installed using code-server [22]), and JupyterLab [50]. All IDE endpoints are password or token protected to prevent malicious actors from accessing user data and code.
Federated Learning servers
The AI4EOSC Platform allows executing federated learning servers [64], allowing AI/ML models to be trained collaboratively across multiple institutions without sharing private data. The platform currently supports the two most widely used and robust frameworks: Flower [10] and NVFLARE [76]. Moreover, we have extended Flower to support token-based client-server authentication [78], with secret management directly implemented via the dashboard. We have also implemented novel aggregation methods, such as FedAvgOpt [79], and incorporated server-side metric differential privacy [77] with fixed clipping, in order to prevent attackers from retrieving training data at inference time from the aggregated model and to defend against client inference attacks.
In order to participate in the federated training, federated clients can be deployed within the AI4EOSC Platform or connect from outside, thus easily accommodating the individual privacy needs of the different data owners. This flexible and distributed approach for training models allows leveraging a wide range of resources in the Cloud continuum that could not support a full ML training cycle by themselves due to their low capacity (e.g., edge devices).
3.1.5 Continuous Integration and Delivery of models
A key component of the AI4EOSC platform, that underpins the MLOps model management is the CI/CD system comprising both platform-enforced and user defined pipelines executed via Jenkins [46]. This system ensures that any AI module integrated in the Marketplace satisfies the same consistent requirements: e.g. by ensuring that published models include the needed metadata, implementing data and model drift detection, or triggering updates on existing services based on a given module.
To achieve this, AI4EOSC maintains a platform-wide pipeline definition that is enforced in all new and existing modules. This pipeline is executed only after the user-defined jobs are executed, and it is optimized in order to execute different tests according to the changes that trigger them, such as metadata updates only or the modification to the source code.
Figure 5: CI/CD pipeline for AI modules
The platform-level CI/CD pipeline is illustrated in Figure 5. As shown, there are a set of steps that comprise software and model quality assurance (e.g., code style, unit testing, security scanning, metadata validation), and a second set of steps that are focused on the delivery of the model in the corresponding components, including keeping track of the steps in the provenance system. From a platform perspective, it ensures the following key aspects:
  • 1.
    Metadata validation: We ensure that the user-defined metadata (4.1) is compliant with the metadata schema that the platform relies on. This is a critical step, taking into account that the PAPI, Dashboard, and the interoperability with other metadata schemes require that accurate and consistent metadata is in place.
  • 2.
    Docker images generation and publication: The platform relies on Docker containers to ensure that the modules are self-contained and executable in a wide variety of systems. Therefore we trigger the build of new images and their publication in the platform registry [^3] https://registry.cloud.ai4eosc.eu/ and Docker Hub [^4] https://hub.docker.com/u/ai4oshub.
  • 3.
    Zenodo release integration: In order to ensure that the code remains accessible, and following the FAIR principles also for software and model assets, we ensure that the Zenodo-GitHub integration is in place. If it is not, we create a deposit and directly upload any pending releases.
  • 4.
    Notification to other platform services: As we will explain in Section 3.2, the platform provides an AI as a Service component that allows to deliver the AI/ML models functionality following a serverless model. In order to let them know that a new version of the model is available, we notify the services via defined web-hooks, providing an updated metadata version, so that they can react to these changes and update the running modules if they are configured to do so. Moreover, we also notify the PAPI, in order to trigger an update of the cached model metadata.
  • 5.
    Notification of provenance updates: The AI4EOSC platform keeps track of the model provenance (3.1.7). Therefore, whenever a model is updated, we notify the provenance system, including the CI/CD gathered metadata (e.g., tests executed), ensuring that the provenance graph always reflects the latest changes.
3.1.6 MLOps components
Extending the DevOps approach to the ML realm, AI4EOSC delivers a series of tools to allow the users to address the key challenges that specifically arise during the ML cycle.
Experiment tracking
Users deploying AI modules (3.1.3) or Development Environments (3.1.4) can log their training metrics into a dedicated MLflow [93] tracking server instance we have deployed, featuring a custom Authentication GUI to be able to integrate it with our authentication system. MLflow credentials, saved in the platform secrets manager and will then be automatically injected in the WMS deployments (3.1.1) created from the Dashboard (3.1.2).
By logging multiple training runs and saving the corresponding model weights with proper versioning, we can enhance the reproducibility of the models. This information is later included in the provenance chain, as we will explain in Section 3.1.7.
Drift monitoring
The AI4EOSC consortium has developed a drift monitoring system, DriftWatch [^5] https://drift-watch.cloud.ai4eosc.eu/. This component allows that whenever AI modules (3.1.3) are created, users can choose to fit a detector on labeled clean and faulty data. Then, at inference time, when the models are deployed in production, the detector can be used to collect and send metrics to the DriftWatch server. From an intuitive UI, users are able to log in and visualize whether drift has occurred, eventually inspecting the data that caused the drift.
In order to detect if a drift is happening or not, a detector is needed. Driftwatch is agnostic to the underlying drift detection framework chosen by the developer such as Alibi-detect [89], Evidently [35], etc. However, AI4EOSC has also developed Frouros [20], a Python library for drift detection in machine learning systems that supports a large combination of classical and more recent algorithms for both concept and data drift detection. Frouros has been integrated within the platform, in order to provide drift detection within AI pipelines easily [86].
3.1.7 Provenance
Provenance is essential for science, and a FAIR-by-design system should take into account asset provenance as one of its pillars. AI4EOSC provides an integrated and extensive provenance system that allows transparent traceability of any AI module (3.1.3), thus improving their reproducibility.
Our system is designed to keep track of all the steps involved in building of a module: from data provenance, to model training, hardware characteristics, etc. Once a module has committed changes to the Git repository, the CI/CD pipeline triggers the fetching of the provenance metadata, gathering information from multiple sources: metadata (4.1) and training resources (3.1.1) from the Platform API (3.1.2), experiment tracking metrics from MLflow (3.1.6) and build information from the CI/CD pipeline itself (3.1.5).
These non-semantic and non-structured metadata fragments are stored in JSON format in a PostgreSQL database [87] managed by an API built with Spring [12]. Then, using the RDF Mapping Language (RML) via CARML [17], we convert the plain JSON metadata into an RDF provenance graph (W3C PROV [90]), optionally using domain specific ontologies. Having the provenance data into RDF allows for an easier conversion to external formats like FAIR4ML [19] or any additional mapping. Moreover, this approach allows us to extend the provenance information with virtually any information that the platform or the user is able to provide to the system.
Figure 6: Example of a provenance graph built for a demo application.
To present the metadata to the users, we provide visual tools integrated in the dashboard (cf. Figure 6), as well as integration with a chatbot assistant (Section 3.3), allowing to perform semantic queries against the graph using natural language.
3.2 AI/ML as a Service system
The AI as a Service (AIaaS), shown in Figure 7, allows user to deploy AI modules (3.1.3) in a serverless inference platform, based on OSCAR [71, 40].
Figure 7: AI4EOSC AI/ML as a Service architecture: container view (C4 notation).
The OSCAR cluster allows to autoscale resources on demand, automatically deploying multiple instances of the model to run predictions as needed. The platform allows both synchronous and asynchronous calls to the endpoints via file uploads to a MinIO object store [60].
The AI4EOSC Dashboard (3.1.2) allows users to automatically create and manage those serverless endpoints for any AI module in the catalog. Thanks to the DEEPaaS API [37] that interfaces the module, users are able to send inference requests to any AI module in a standardized way.
Leveraging on those serverless endpoints, the AI4EOSC Platform allows to create AI pipelines where users can compose multiple inference requests to different AI models, as well as data preprocessing and postprocessing steps.
Pipelines can currently be created with two different visual composition tools: Node-Red [65] whose instances are managed via the OSCAR cluster, and Elyra [29], integrated with EGI Jupyter Notebooks. Users can choose between these two tools to compose their pipelines based on their working environment needs and prior experience.
3.3 Large Language Models Services system
This system, described in Figure 8, is designed to provide a federated inference system focused on the needs of Large Language Models (LLMs). The system leverages vLLM [53] and LiteLLM [55] in order to provide an easy way to include additional LLM providers, that can be geographically distributed. Moreover, it is integrated with the platform secret management system, in order to be able to manage API keys for the users. This system can be directly used via its OpenAI compatible API endpoints, in order to be integrated with any user tool, or via a conversational engine. Moreover, this system is key to provide AI4EOSC platform users with an specific support agent, able to support users with direct answers from the documentation knowledge base.
Figure 8: AI4EOSC Large Language Model Services architecture: container view (C4 notation).
3.4 Platform orchestration system
This system, depicted in Figure 9 is in charge of assisting resource providers and platform operators by automatically provisioning the resources needed to deploy the rest of the platform services. In the AI4EOSC platform, we use the INDIGO-PaaS Orchestration system [26, 80] to handle the deployment of platform components across the different IaaS platforms made available by the federated Cloud providers.
Figure 9: AI4EOSC platform orchestration architecture: container view (C4 notation).
The system is composed of different micro-services:
  • 1.
    The Infrastructure Manager [15, 16], who is responsible for the submission of the deployment to the proper Cloud provider. The provider is selected according to the AI-Ranker [38], which generates a ranked list of providers by combining the prediction of deployment success/failure with the estimated deployment creation time.
  • 2.
    The Federation-registry [81], which is the configuration management database for dynamic cloud federation. Built using FastAPI [73] and neo4j [62], the Federation-registry is able to efficiently manage the federation-related information and SLA agreements.
  • 3.
    The INDIGO-PaaS Orchestrator Dashboard, which provides a user-friendly interface for managing and monitoring the TOSCA-based deployments, built on Flask [68] and MySQL [61].
With the developed TOSCA [69] templates, the PaaS Orchestration layer is able to automate the deployment of the platform components, including the Workload Management System (3.1.1) and the AI as a Service platform (3.2).
These strong reproducible recipes make the AI4EOSC platform a suitable solution to be easily adopted by external communities wishing to profit from the technologies developed here. Moreover, the adoption of well known technology standards allow the Cloud-enabled resource providers to easily join the federation and contribute to the existing AI4EOSC platform.
4 Other platform design considerations
4.1 AI4EOSC metadata
In the current data-driven ecosystem, rich metadata is of paramount importance in order to comprehensively describe a given asset in an homogeneous way. Well defined metadata allows also to work towards interoperability, both at the platform and at the AI/ML model level. To ensure that all the AI modules developed in the platform provide consistent metadata, the AI4EOSC Platform enforces that each AI module (3.1.3) or tool (3.1.4) that is registered in the AI4EOSC catalog must include a valid metadata file following a defined schema implemented via JSON Schema [47].
The AI4EOSC metadata schema includes both user-defined fields (some of them optional) as well as fields automatically filled from external sources like License and modification dates from GitHub. User defined-fields include title, summary, description, DOI, external links (e.g. to dataset or model weights) and tags such as the software libraries used or the data types used as input.
Once integrated, the Platform API retrieves the module’s metadata and serializes it in different formats (JSON-LD, RDF Turtle), being able also to transform it into different application profiles, such as the MLDCAT Application Profile [84]. This allows for a greater interoperability of our models with external catalogs, allowing semantic interoperability with formats that are compatible with MLDCAT-AP.
4.2 Platform extensibility and interoperability
One of the design decisions of AI4EOSC is that any scientific platform must be able to adapt to the needs of their user community, and not viceversa. In order to be able to comply with this design motto, our systems are built with extensibility in mind, meaning that we are able to integrate, to the technical possible extent, with external tools easily. In this section we present the three AI4EOSC extensibility and interoperability axes.
4.2.1 External deployment platforms
All executable assets in the platform are based on Docker containers, that are being released via reproducible builds into a public Docker Hub organization [^6] https://hub.docker.com/u/ai4oshub. This means that the AI modules (3.1.3) are easily deployable at external cloud providers and in any other infrastructure where Docker is available. From the Dashboard (3.1.2), we give users direct access to deploy in any external cloud provider via the Infrastructure Manager [32], that is part of our orchestration system (3.4). We have also developed connectors to be able to deploy on the EOSC EU Node [30], allowing users to leverage its resources.
4.2.2 External AI catalogs
Due to our AI modules design (3.1.3), which hides the module internal workings behind a unified API [37], we are able to easily build connectors with external AI module catalogs. These connectors allow to deploy those external modules in our platforms. To demonstrate this, we have built a connector to the BioImage Model Zoo catalog [66], that leverages their standardized metadata to map their functionality into our platform, and making them available directly in the Dashboard (3.1.2). This feature allows the allowing one-click deployments of those external modules, allowing the BioImage Model Zoo users to continue working with their tools, but making their outcomes available in the AI4EOSC marketplace.
4.2.3 External datasets
AI/ML models rely heavily on data for development and training. Therefore, providing users with the ability to directly download data is an important feature. Additionally, in a metadata-rich platform like AI4EOSC, users can specify which dataset has been used to train a given model by referencing their DOI or URL. This capability is particularly valuable because it makes it easy to reproduce results, ensuring transparency and reliability in the research process. Our dashboard enables users to designate the data they wish to synchronize at launch time, streamlining the data management process.
We implement this functionality transparently via sidecar tasks (3.1.1) and extending the the DataHugger tool [14] with additional integrations, currently including Zenodo [33], Data Europa [34], Dryad [28] or SeaNoe [82].
4.2.4 External storages
In addition to the platform storage, we are able to connect user deployments with any external storage, including Nextcloud, S3 [5], Owncloud [67], Dropbox [27], Google Drive [3], etc. In order to do so, we allow users to link the platform with the supported storage systems (e.g., NextCloud) or manually providing their user credentials for other storage providers.
5 Use cases and usage scenarios
5.1 Common usage scenarios
In this section we will explain how scientists can leverage the AI4EOSC Platform all along the ML lifecycle.
5.1.1 Developing a model
Developing an AI/ML model in AI4EOSC begins with a reproducible, MLOps-driven workflow that ensures traceability and compliance with FAIR principles. Researchers and platform users can deploy a pre-configured development environment (Codespace) via the Dashboard, selecting their preferred IDE (VSCode or JupyterLab) and utilizing their preferred framework (e.g., TensorFlow, PyTorch). The environment is linked to version-controlled repositories (GitHub) and experiment tracking (MLflow) tools, enabling users to capture metadata into it. The repository structure is set up based on one of the AI4EOSC software templates.
As already mentioned, the platform enforces a structured JSON schema for model metadata (e.g., dependencies, datasets, licenses), ensuring interoperability with external catalogs (MLDCAT-AP). Every code commit triggers the CI/CD pipeline, that ensures that the development steps are logged into the W3C PROV-compliant provenance graph, enabling full lineage tracking. Moreover the same pipeline, enforced at platform level, builds the Docker image and refreshes all the relevant components, so that the new module is ready to be used.
5.1.2 Training a model
Training in AI4EOSC leverages our federated Workload Management System (WMS) to dynamically allocate resources (GPUs and CPUs) across distributed providers, resource allocation, energy consumption and the overall performance. Researchers configure training jobs via the Dashboard, specifying the data sources that need to be linked with their deployment (e.g., Zenodo datasets), using data annotation tools within the platform, or utilizing federated learning by leveraging the integrated federated learning server.
The generated trained model can then be integrated in the module catalog, as in the model development case (5.1.1), performing the same MLOps and CI/CD approach previously described.
5.1.3 Deploying a module
A deployment in AI4EOSC is provided via different options, according to the user and application needs, and the resources involved. AI4EOSC covers the computing continuum in a multi-faceted approach:
  • 1.
    Permanent deployment within the WMS, if they want their module to be permanently deployed for low-latency or prototyping.
  • 2.
    External platforms with one-click deployment to external or commercial clouds (AWS, GCP) via the Infrastructure Manager [16]
  • 3.
    On-demand inference using the AI as a Service component (3.2) based on OSCAR, following a serverless MLOps-optimized approach with on-demand autoscaling.
  • 4.
    Edge/Cloud Continuum: Models can be deployed to HPC clusters or edge devices (Raspberry Pi [75]).
5.2 Platform adoption
Besides building a generic and horizontal platform, and in order to foster the adoption of the technology, we have closely worked with several initiatives in order to build a customized version the platform, tailored to the user community needs. The interactions with those communities has led to a virtuous circle, where the platform has guided them on their adoption of AI and they have in turn provided feedback for a user-driven platform development.
5.2.1 iMagine
The iMagine [^7] https://www.imagine-ai.eu/ initiative is a European Union–funded research project that focuses on developing and operating an AI-enabled imaging platform for aquatic science. It aims to give researchers working on oceans, seas, coastal and inland waters access to rich image datasets, advanced AI-powered analysis tools and best-practice guidance for scientific image processing. The iMagine platform is built on the software stack developed within AI4EOSC project.
During iMagine, the Competence Center team provided support to the usersa and collaborated closely with the developers of AI4EOSC project, providing feedback to enhance the platform. This partnership enabled hands-on experience across the full AI workflow, from image collection and dataset preparation to model training and deployment for aquatic science applications, ultimately resulting in the publication [8].
5.2.2 KMD4EOSC
KMD4EOSC [^8] https://kmd4eosc.pl/ is a Poland national project for data storage and sharing, and efficient processing of large volumes of data in HPC, BigData, and AI Models. The project is co-funded by European Funds for a Modern Economy Program and the European Union. The project is one of the flagship projects of the Polish Research Roadmap.
The KMD4EOSC project designs, builds and implements infrastructure, services and applications for storing and collecting scientific and economic data, as well as for processing and analyzing them and making them available to the scientific community and the economy in an efficient manner and in accordance with good practices, requirements and regulations. AI4EOSC has been deployed within the project to serve as a platform for providing the whole lifecycle of AI/ML models, integrated with repositories, datasets and lower infrastructure and data management services.
5.2.3 AI4Life
AI4Life [^9] https://ai4life.eurobioimaging.eu/, a Horizon Europe funded project to empower life science researchers to harness the full potential of AI/ML methods for bioimage analysis, began collaborating with AI4EOSC in order to leverage the AI4EOSC software stack and platform. The AI4EOSC and AI4Life teams worked on both platforms integration, as a means to implement and pilot a model for interoperability across research infrastructures, aligned with FAIR principles. This integration enables efficient sharing and use of AI/ML models and computing resources across domains.
Central to this effort is deploying models from the BioImage Model Zoo within the AI4Life cloud marketplace. Compatibility was ensured through defined execution environments such as PyTorch, standardized metadata rich input and output formats, and consistent prediction pipelines, allowing deployment via Dockerfile and DEEPaaS compatible APIs.
6 Conclusion and Future Work
With a fragmented AI ecosystem increasingly dominated by large corporations, a scalable open-source solution putting the focus on reproducibility and interoperability is needed. To address this challenge, in the context of science and research, we have here presented the AI4EOSC platform. We provide a comprehensive set of tools that target the whole AI/ML lifecycle, backed with easy to use interfaces and dashboards, a federated workload management system to train models, options to deploy inference models along the whole Cloud Continuum, the ability to create composable inference workflows paired with our AI/ML as a Service serverless framework, built-in AI model provenance to foster trustworthiness, and an extensive CI/CD pipeline that ensures AI code quality and enables MLOps practices.
The AI4EOSC architecture provides a modular and extensible reference that allows to build upon it, enabling the integration of external services and catalogs. It provides an overall customization system that has been adopted by relevant external research communities, and that is enabling and advancing FAIR development of AI/ML in science and within the EOSC. In these lines, we are working with specific user communities in order to integrate their community tools (e.g., data repositories, virtual research environments) with the platform.
Concerning future work, we plan to keep evolving the platform along several axes. Specifically, we plan to perform extensive work in interoperability, and namely we will extend the semantic support of ML assets, like datasets or models, to be able to seamlessly import, export and connect them in the platform, fostering semantic interoperability following open standards such as Croissanl-ML [2]. In addition, we plan to extend the usage of the Experiment Tracking component to serve as a model registry, enriching the models with semantic information. We will also further strengthen component integration, for example, by allowing the use of any existing AI4EOSC AI module as an annotation assistant inside the CVAT tool, using the AIaaS serverless endpoints.
Besides semantic interoperability, we plan to work on the technical interoperability. More specifically, we plan to implement the OAI-PMH protocol at the marketplace level, making it possible to automatically harvest the module metadata. Additional integrations and interoperability with external marketplaces and model repositories (both science related and generalistic) are in our roadmap.
Lastly, we are also working towards making the platform aware of the energy consumption, both at the control plane and the user workloads. This will allow us to calculate the carbon and water footprint, integrating fine-grained and accurate impact measurements from Wattnet [18], and exposing them to the users, via the module metadata as well as upon workload submission.
7 Acknowledgments
The authors acknowledge Marcos Lloret Iglesias, the AI4EOSC project manager, for his constant support and help towards bringing the project to fruition. The authors acknowledge the funding and support from the AI4EOSC project “Artificial Intelligence for the European Open Science Cloud”, that has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement number 101058593. ALG also acknowledges the support from the Consejería de Educación, Formación profesional y Universidades of the Gobierno de Cantabria via the “Actividad estructural para el desarrollo de la investigación del Instituto de Física de Cantabria” project.